OpenStudy (anonymous):

For Me To Pass Me Exam help!! Thankss : A public library has fiction and nonfiction books in the ratio three colon two. If the library has three hundred sixty fiction books, how many nonfiction books does it have?

OpenStudy (anonymous):

if ratio is 3:2 then we can divide by the total: 5 and then multiply by 2 to get the number of nonfictions 360/5 = 72 72*2 = 144

OpenStudy (anonymous):

another way to think about it is for some number x fiction = 3x non fiction = 2x and total = 3x + 2x = 5x = 360 5x = 360 x = 72 nonfiction = 2x = 2*72 = 144

OpenStudy (anonymous):

Thanks..

OpenStudy (anonymous):

is it ok? or have i made it confusing?

OpenStudy (anonymous):

it's ok

OpenStudy (anonymous):

good :)